OpenStudy (anonymous):

Find the range of the given data set. 5/8, 3/4, 1 1/8, 1 1/4 1 1/8?

OpenStudy (anonymous):

largest minus smallest = range

OpenStudy (anonymous):

3/4?

OpenStudy (anonymous):

@telliott99 ?

OpenStudy (anonymous):

11/4-5/8

OpenStudy (anonymous):

17/8

OpenStudy (anonymous):

5/8... ?

OpenStudy (anonymous):

1 1/4 = 10/8 10/8 - 5/8 =?

OpenStudy (anonymous):

Good.